Hello, can someone help with this please? Q Search Sheet Σ AutoSum ΑΘ, Z Insert Delete Format Conditional Format Formatting as Table Cell Styles Clear Sort & Filter DWorkshop 1 Problems Student Q Search Sheet Review View Σ AutoSum ΑΘ, Wrap Text Number Z DOO Home Insert Page Layout Formulas
Answer #1

HOME INSERT FILE PAGE LAYOUT FORMULAS DATA REVIEW VIEW A X Cut - 11 E Wrap Text Calibri General Ee Copy - Paste E E E Merge &Excel formulas

X Cut E Copy Format Painter E AutoSum - - |11 - A A= Wrap Text Calibri General U Fill- Fill- Conditional Format as Cell Forma

  • Step 1: Ending retained earnings = $60,000 given in 'D29'. Input this to 'C48'
  • Step 2: Interest payable = Total liabilities - Loan payable = $1,000
  • Step 3: Find capital balance. Capital stock = Total liabilities and stockholders' equity - (total liabilities + retained earnings) = $228,000
  • Step 4: Find total assets. Total assets = Total liabilities
  • Step 5: Find Cash. Cash = Total assets - (accounts receivable+land) = $54,000
  • Step 6: Beginning retained earnings + Net income - Dividends = Ending retained earnings.
    • i.e. Beginning retained earnings + Net income = Ending retained earnings + Dividends = 60000+30000 = $90,000
    • Now, Beginning retained earnings + Net income = $90,000
    • Net income = $90,000 -  Beginning retained earnings = $90,000 - $30,000 = $60,000
  • Step 7: Net income in the income statement = 'D26' i.e. as found above
  • Step 8: Revenue - Expenses = Net income
    • Revenue - (Wages+Interest) = Net income
    • Wages = Revenue - (Interest+Net income) = $200,000 - ($3,000+$60,000) = $137,000